A rectangle has an area of 36 square meters. Its length is 2square-root of 3 meters. In exact form, what is the permieter of the rectangle?
A right triangle has a 24-cm perimeter, and its hypotenuse is twice as long as its shoter leg. To the nearest tenth of a cm, fnd the lengths of all three sides of this triangle.

Answer by nycsharkman(136)

You have two questions here.
Question(1).
for the long side of the rectangle is x,
the short side of the rectangle is y,
x*y=36m^2
x=2sqrt3 y=36/2sqrt3=6sqrt3
perimeter=2(x+y)=2(2sqrt3+6sqrt3)=16sq...

Question (2)
short leg of triangle=a
long leg of triangle=b
hypotenus=h=2a
h^2=a^2+b^2
b^2=3a^2
b=asqrt3
h+a+b=3a+b=24
3a+asqrt3=24, a(3+sqrt3)=24, a=24/(3+sqrt3)
if we multiply the denominator with (3-3sqrt3)
then a=(12-4sqrt3)
b=sqrt3*a=sqrt3*(12-4sqrt3)=12sqrt3-12
h=2a=24-8sqrt3